How many grams of oxygen gas must react to give 2.90 g of ZnO?
2Zn(s)+O2(g)→2ZnO(s)
Express your answer with the appropriate units.
1
Expert's answer
2020-11-20T08:07:49-0500

n=m/M

M (ZnO)=81.38 g/mol

n (ZnO)=2.90/81.38=0.04 mol

n(O2)=1/2 n (ZnO)=0.04/2=0.02 mol

M(O2)=16 g/mol

m (O2)=0.02 x 16 = 0.32 g